# Towards Inclusive ASR: Investigating Voice Conversion for Dysarthric Speech Recognition in Low-Resource Languages

### v2

@@ 1 Introduction
A major challenge in improving ASR for dysarthric speech is the severe scarcity of annotated dysarthric speech datasets [4, 5]. Collecting such data is inherently difficult, as recording sessions can be physically demanding for individuals with dysarthria, resulting in limited availability of large, high-quality corpora. Currently, there are only around ten publicly available dysarthric speech datasets [6], with four of them covering English [7, 8, 9, 3]. The situation is even more critical for non-English languages, where dysarthric speech resources are largely nonexistent. This severe data imbalance restricts the development of robust ASR models for non-English dysarthric speech, exacerbating accessibility challenges for individuals with dysarthria across diverse linguistic communities.
@@ 3.1 Datasets
PC-GITA [28] includes 50 patients with PD and 50 healthy subjects evenly matched in age and gender, all native Spanish speakers from Colombia. Each participant read 10 phonetically balanced sentences and 25 isolated words representing the phonetic inventory of Colombian Spanish.
@@ 5.1 Objective Evaluation: Classification
We train an XGBoost classifier [38] for each language using prosodic features extracted with DisVoice [39] to distinguish between dysarthric and healthy speech. The model is trained on dysarthria datasets using an 8:2 group-stratified train-test split. We optimize the classifier through grid search and select the best-performing model. The selected classifier is then used to evaluate whether the generated speech is perceived as healthy or dysarthric3. The F1-scores on the test set are as follows: 69.98% (Spanish-All), 82.16% (Italian-All), 93.99% (Tamil-All).
@@ 5.2 Subjective Evaluation: Perceptual Tests
For the subjective evaluations, we selected 10 audio samples from each group: healthy and dysarthric speech from the dysarthric datasets, and generated speech from each augmentation method. The samples were balanced across severity levels and randomly selected within each. For each language, two native speakers with no prior training in speech pathology provided judgments in two tasks. We provided information on perceptual characteristics of dysarthric speech before evaluation3. The first evaluation was conducted on dysarthric datasets, where participants rated whether the audio sounded dysarthric (“Dys.?”) using a 1-to-4 Likert scale [40], ranging from “definitely healthy” to “definitely dysarthric.” As the second evaluation, we asked the participants to apply the same scale to generated speech samples, asking how closely generated speech approximates the dysarthric or healthy speech characteristics. For the generated samples, we asked additional question on the naturalness in the target language (“Your Lang.?”). 1 to 4 Likert scale was used, ranging from “Not at all my language” to “Definitely my language.” This helps to provide information on how well the linguistic characteristics were preserved by the augmentation methods.

### v3

@@ 1 Introduction
A major challenge in improving ASR for dysarthric speech is the scarcity of annotated dysarthric speech datasets [4, 5]. Collecting such data is inherently difficult, as recording sessions can be physically demanding for individuals with dysarthria, resulting in limited availability of large, high-quality corpora. Currently, there are only around ten publicly available dysarthric speech datasets [6], with four of them covering English [7, 8, 9, 3]. The situation is even more critical for non-English languages, where dysarthric speech resources are largely nonexistent. This severe data imbalance restricts the development of robust ASR models for non-English dysarthric speech, exacerbating accessibility challenges for individuals with dysarthria across diverse linguistic communities.
@@ 3.1 Datasets
PC-GITA [28] includes 50 patients with Parkinson’s Disease (PD) and 50 healthy subjects evenly matched in age and gender, all native Spanish speakers from Colombia. Each participant read 10 phonetically balanced sentences and 25 isolated words representing the phonetic inventory of Colombian Spanish.
@@ 5.1 Objective Evaluation: Classification
We train an XGBoost classifier [38] for each language using prosodic features extracted with DisVoice [39] to distinguish between dysarthric and healthy speech. The model is trained on dysarthria datasets using an 8:2 group-stratified train-test split. We optimize the classifier through grid search and select the best-performing model. The selected classifier is then used to evaluate whether the generated speech is perceived as healthy or dysarthric1. The F1-scores on the test set are as follows: 69.98% (Spanish-All), 82.16% (Italian-All), 93.99% (Tamil-All).
@@ 5.2 Subjective Evaluation: Perceptual Tests
For the subjective evaluations, we selected 10 audio samples from each group: healthy and dysarthric speech from the dysarthric datasets, and generated speech from each augmentation method. The samples were balanced across severity levels and randomly selected within each. For each language, two native speakers with no prior training in speech pathology provided judgments in two tasks. We provided information on perceptual characteristics of dysarthric speech before evaluation1. The first evaluation was conducted on dysarthric datasets, where participants rated whether the audio sounded dysarthric (“Dys.?”) using a 1-to-4 Likert scale [40], ranging from “definitely healthy” to “definitely dysarthric.” As the second evaluation, we asked the participants to apply the same scale to generated speech samples, asking how closely generated speech approximates the dysarthric or healthy speech characteristics. For the generated samples, we asked additional question on the naturalness in the target language (“Your Lang.?”). 1-to-4 Likert scale was used, ranging from “Not at all my language” to “Definitely my language.” This helps to provide information on how well the linguistic characteristics were preserved by the augmentation methods.